OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (precal):

multiply -4 to the second equation

OpenStudy (precal):

-4(x-y=3) -4x+4y=-12

OpenStudy (precal):

add that equation to the first and solve for y. then sub y into the second equation to find x

Directrix (directrix):

x = 3